# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

### v4

@@ 6.1.2 Supply disruption
Supply disruption can be modeled as a proportional reduction in both the capacity and the maximum vehicle accumulation on the MFD profile, as represented in Fig. 7(b). As discussed in Section 5.1, only the MFD profile of the inner region is reduced as a critical scenario in our simulation, where the outer region remains intact. Figure 11 presents the performance curves of the studied algorithms under supply disruption. In addition to MPC (orange), baseline RL (gray), and the proposed antifragile RL (blue), the MPC-MHE variant (brown) is also included. By modeling the supply disruption magnitude coefficient $r$ as an unknown system parameter, the MHE module estimates $r$ in real time to capture the extent of the MFD reduction. Fig. 10(a) shows the performance curves of the four studied algorithms under the no supply disruption for the first $50$ episodes, followed by linearly declining MFD during episode $50-75$. Unlike the performance under demand disruption in Fig. 8(a), where MPC performs better at lower disruption magnitudes, the TTS curve of the proposed antifragile RL algorithm remains consistently similar or below those of the other methods, indicating superior performance across the entire range of disruption.
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 10: Performance curves under incremental supply disruptions
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 11: Performance curves under incremental supply disruptions with uncertainty
@@
The performance gain is quantitatively analyzed in Fig. 10(b), which shows an average TTS reduction of $16.7\%$ from the proposed antifragile RL compared to the baseline RL algorithm under incremental supply disruption, and reaches an ultimate performance improvement of $39.4\%$. It is worth noting that the MPC-MHE curve in brown also delivers strong performance, achieving a $31.4\%$ TTS reduction at episode $75$, highlighting the effectiveness of incorporating the supply disruption magnitude coefficient $r$ to be estimated within the MHE framework. Still, the performance does not quite match that of the antifragile RL algorithm under high magnitudes of supply disruption. Moreover, MHE relies on accurately modeling both the system dynamics and the evolution of the model itself, which is a highly challenging requirement in real-world scenarios. Another shortcoming of MHE is the marginally worse performance when no supply disruption is present, i.e., during episode $0-50$, showing a $4.0\%$ reduction compared to MPC.
@@
The distribution skewness up to the $n$-th episode under disruption is demonstrated in Fig. 10(c). The proposed antifragile RL method exhibits noticeably lower skewness across episodes $50-75$ compared to the other approaches, with a final skewness of $0.49$, whereas the skewness of the other three methods hovers around $1$. As discussed in Section 6.1.1, the spike in the skewness curve of the antifragile RL algorithm around episode $60-65$ may be caused due to the mixture of training data under both disrupted and undisrupted scenarios. Note that the baseline RL algorithm exhibits a plateau in skewness toward the end of the simulation, indicating the network has already succumbed to a gridlock. When a network is gridlocked, TTS increases linearly with the current vehicle accumulation. This implies that beyond a certain threshold, demand disruption will only lead to a linear TTS growth instead of exponential, while supply disruptions primarily accelerate the onset of gridlock without significantly affecting the eventual TTS. As a result, when overreaching the disruption threshold, the skewness of TTS initially plateaus and then gradually approaches zero. Such a phenomenon is more pronounced in the below Fig. 11(c). Since the skewness value being zero suggests the system is neither fragile nor antifragile, the proposed algorithm is applicable under typical disruptions through antifragile control, while not designed to handle catastrophic scenarios where the network rapidly descends into gridlock.
@@
When supply disruption uncertainty is introduced into the simulation, as shown in Fig. 11, the general trend is similar to that observed under demand disruption with uncertainty. The performance curves of MPC, MPC-MHE, and the antifragile RL algorithm all deteriorate in the presence of stochasticity, while the baseline RL algorithm surprisingly shows improved performance. As discussed earlier, this improvement can be attributed to the onset of gridlock and its effect on supply disruptions. When stochasticity amplifies the supply disruption magnitude coefficient, its impact is minimal since the network is already gridlocked. In contrast, when stochasticity attenuates the coefficient, it reduces TTS and enhances performance. As a result, the overall performance of the baseline RL algorithm appears to improve. Nevertheless, the performance gain of the proposed algorithm is demonstrated quantitatively in Fig. 11(b), showing an average improvement of $14.6\%$ and reaching $27.1\%$ at the final episode, while MPC and MPC-MHE show an ultimate $6.4\%$ and $19.1\%$ performance gain, respectively. 11(c) summarizes the skewness of the studied algorithms, with our proposed antifragile RL achieving the lowest value of $0.70$. Although the ultimate distribution skewness of the baseline RL reaches $0.79\%$, its peak skewness is $0.90\%$ at episode $71$.
@@ 6.2 Performance under real-world limited observability
As discussed in Section 4.1, obtaining real-world measurements of $n_{ij}$ and $q_{ij}$ is highly challenging. Therefore, based on Eq. 8 and 9, we evaluate the performance of the baseline and proposed antifragile RL-based algorithms under conditions of real-world limited observability. It is important to highlight that the performance of the proposed algorithm under limited observability is not directly comparable to MPC or MPC-MHE under full observability, given their fundamentally different observability requirements. Fig. 13 and 13 illustrate the performance of algorithms in the presence of linearly increasing demand disruptions, modeled either deterministically or with disruption magnitude uncertainties.
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 12: Performance curves under incremental demand disruptions
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 13: Performance curves under incremental demand disruptions with uncertainty
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 14: Performance curves under incremental supply disruptions
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 15: Performance curves under incremental supply disruptions with uncertainty
@@
In Fig. 12(a), while the TTS curve of the baseline RL algorithm in light coral is approximately the same under both full and limited observability, the performance curve of the antifragile RL algorithm under limited observability in light teal falls between that of the baseline RL and the antifragile RL with full observability. This indicates that the algorithm sacrifices some performance in exchange for depending on more realistic observable measurements. The TTS difference in Fig. 12(b) marks the performance difference between the baseline and antifragile RL-based algorithms under limited observability, taking an average of $6.1\%$ performance improvement and reaching $15.0\%$ at the end of the simulation. It should be highlighted that the performance trade-off to achieve realistic observability, i.e., the performance difference between limited (light teal) and full observability (blue) of the proposed algorithm, is an average of $4.0\%$. The distribution skewness is quantitatively illustrated in Fig. 12(c). Although taking a higher skewness value at the onset of disruption, the proposed method yields a rather low ultimate skewness of $0.49$ under limited observability, lying between the skewness observed with full observability ($0.46$) and that of the baseline RL under limited observability ($0.62$).
@@
When under demand disruption with magnitude stochasticity as shown in Fig. 13(a), the proposed RL algorithm suffers from slightly larger deterioration compared to the baseline RL, mainly because the performance of the baseline RL is close to network gridlock. Still, the TTS difference in Fig. 13(b) showcases the superiority of the proposed algorithm with an average of $4.2\%$ performance gain and reaches $8.2\%$ at the final episode. The performance loss for the trade-off of observability is $4.3\%$. The skewness under disruption stochasticity exhibits some volatility, which shows a value of $0.61$ and is identical to that of the baseline RL under limited observability.
@@
In Fig. 15 and 15, the performance of both algorithms under supply disruptions with limited observability follows similar trends to those under demand disruption. The shaded area in Fig. 13(b) exhibits the TTS difference between the two algorithms with limited observability, which yields an average performance gain of $11.1\%$ and achieves a final value of $30.7\%$. The observability trade-off is $5.5\%$, taking an even smaller percentage of the average gain than under demand disruption. Likewise, its skewness $0.70$ at episode $25$ is also positioned between the baseline RL with limited observability and the antifragile RL with full observability. When under disruption magnitude stochasticity in Fig. 15(a), as explained, the performance of the baseline RL method improves at high disruption magnitude, because the network is close to gridlock. When stochasticity brings in higher and lower magnitudes of disruptions, stronger disruption has little effect on the performance, whereas weaker disruption leads to better performance. An intuitive proof can be observed at episode 70, where the TTS of the baseline RL both with and without uncertainty in Fig. 14(a) and Fig. 15(a) is around $1\cdot 10^{8}$ seconds, indicating the gridlock taking place shortly afterwards. Fig. 15(b) exhibits an ultimate performance gain of $12.5\%$. Although this scenario shows the highest observability trade-off of $10.1\%$, the skewness of the proposed antifragile RL is rather low with a value of $0.69$ compared to the other algorithms.

### v5

@@ 6.1.2 Supply disruption
Supply disruption can be modeled as a proportional reduction in both the capacity and the maximum vehicle accumulation on the MFD profile, as represented in Fig. 7(b). As discussed in Section 5.1, only the MFD of the inner region is reduced as a critical scenario in our simulation, where the outer region remains intact. Figure LABEL:fig:_s presents the performance curves of the studied algorithms under supply disruption. In addition to MPC (orange), baseline RL (gray), and the proposed antifragile RL (blue), the MPC-MHE variant (brown) is also included. By modeling the supply disruption magnitude coefficient $r$ as an unknown system state variable, the MHE module estimates $r$ in real time to capture the extent of the MFD reduction. Fig. LABEL:fig:_incr_s shows the performance curves of the four studied algorithms under the no disruption for the first $50$ episodes, followed by linearly reduced MFD during episode 507550-75.
@@
The performance gain is further quantitatively analyzed in Fig. LABEL:fig:_incr_s_p, which shows an average TTS reduction of $16.7\%$ from the proposed antifragile RL compared to the baseline RL algorithm under incremental supply disruption, and reaches an ultimate performance improvement of $39.4\%$. It is worth noting that the MPC-MHE curve in brown also delivers strong performance, achieving a $31.4\%$ TTS reduction at peak disruption, highlighting the effectiveness of incorporating the supply disruption magnitude coefficient $r$ to be estimated within the MHE framework. Still, MHE relies on accurately modeling both the system dynamics and the evolution of the model reduction, which is highly challenging in real-world scenarios.
@@
The distribution skewness up to the $n$-th episode under disruption is demonstrated in Fig. LABEL:fig:_incr_s_sk. The proposed antifragile RL method exhibits noticeably lower skewness across episodes 507550-75 compared to the other approaches, with a final skewness of $0.49$, whereas the skewness of the other three methods hovers around $1$. Note that the baseline RL algorithm exhibits a plateau in skewness toward the end of the simulation, which is caused by the fact that the network has already succumbed to a gridlock. And when the network is gridlocked, TTS merely increases linearly instead of exponentially with the growing vehicle accumulation. As a result, when overreaching the disruption threshold, the skewness of TTS initially plateaus and then gradually approaches zero. Such a phenomenon is more pronounced in the following Fig. LABEL:fig:_incr_s_sk_un, indicating that the proposed algorithm is only applicable under typical disruptions, while not for handling catastrophic scenarios where the network rapidly descends into gridlock.
@@
When supply disruption uncertainty is introduced into the simulation, as shown in Fig. LABEL:fig:_s_un, the general trend is similar to that observed under demand disruption with uncertainty. The performance curves of MPC, MPC-MHE, and the antifragile RL algorithm all deteriorate to some extent in the presence of stochasticity. Interestingly, the performance of the baseline RL algorithm appears to improve at peak disruption compared to Fig. LABEL:fig:_incr_s. Such improvement can be attributed to the onset of gridlock and its effect on supply disruptions. When stochasticity amplifies the supply disruption magnitude coefficient, its impact is minimal since the network is already gridlocked. In contrast, when stochasticity leads to an attenuated coefficient, it reduces TTS and enhances performance. Nevertheless, the performance gain of the proposed algorithm is demonstrated quantitatively in Fig. LABEL:fig:_incr_s_p_un, showing an average improvement of $14.6\%$ and reaching $27.1\%$ at the final episode, while MPC and MPC-MHE show an ultimate $6.4\%$ and $19.1\%$ performance gain. LABEL:fig:_incr_s_sk_un summarizes the skewness of the studied algorithms, with our proposed antifragile RL achieving the lowest value of $0.70$. Although the ultimate distribution skewness of the baseline RL reaches $0.79\%$, its peak skewness is $0.90\%$ at episode $71$.
6.1.3 RL agent decision-making and explainability analysis
Previous sections report performance across the full 75 training episodes for both the state-of-the-art and the proposed antifragile algorithms, and here we examine individual representative episodes to address the explainability of the two RL-based algorithms. Specifically, we present the perimeter control variables $u_{ij}$ yielded from the RL agents and the consequent vehicle accumulation $n_{i}$.
Fig. 8 highlights episodes 50 and 75, which are the final episode before any disruption is introduced and the final episode with the highest disruption magnitude, respectively. The upper plot in Fig. 8(a) shows $u_{12}$ of both the baseline RL agent in gray and the proposed algorithm in blue, with the shaded area representing one standard deviation. Note that since the outer region is much less susceptible to traffic, both agents learn $u_{21}$ fairly easily to be at the upper bound $0.9$. The peak demand occurs at $t=1800$, resulting in the lowest $u_{12}$ for the baseline algorithm in gray. However, $n_{2}$ from the baseline algorithm depicted as the dashed gray curve in the lower plot has clearly exceeded the critical accumulation $n_{2,crit}$ shown as the red dashed line, indicating the sub-optimal decision-making of $u_{12}$. Clearly, the learned policy of the baseline algorithm with state defined as $s_{k}=[n_{ij,k},q_{ij,k}]$ is heavily influenced by $q_{ij,k}$, since $u_{12}$ from the baseline is nearly symmetric to the peak demand as in Fig. 7(a). It should also be noted that, since MFD is largely a concave function, a small deviation between the proposed $n_{2}$ as the blue dashed curve and $n_{2,crit}$ will only lead to very minor performance loss, while an increasing deviation causes exponential loss.
(a) Episode 50 no disruption.
(b) Episode 75 demand disruption.
(c) Episode 75 supply disruption.
Figure 8: Representative episodes for the perimeter control variables and the vehicle accumulation under the last episode of no disruption, peak demand disruption, and peak supply disruption scenarios.
Fig. 8(b) demonstrates $u_{ij}$ and $n_{i}$ of the final episode $75$ as the demand disruption magnitude gradually intensifies over episode 507550-75. The solid and dashed horizontal red lines represent the critical accumulation for either region $n_{1,crit}$ or $n_{2,crit}$, respectively. Due to the demand disruption on top of the base demand, both algorithms reach the lower bound $0.1$ some time before $t=1800$. The proposed RL agent remains approximately at the lower bound for a longer period, whereas the baseline RL agent quickly rebounces as a reaction to the waning of peak demand after $t=1800$, resulting in exceedingly high $n_{2}$ represented by the dashed gray curve. In contrast to Fig. 8(b), Fig. 8(c) shows that the baseline $u_{12}$ in gray has not even reached the lower bound and rebounces only shortly after the peak of demand $t=1800$, while the $u_{12}$ of the proposed algorithm in blue remains at the lower bound first and then manages the accumulation in the inner region as the dahsed blue curve somewhere around $n_{1}=3000$ after $t=5400$, corresponding to the compromised maximal completion rate under the peak supply disruption in Fig. 7(b). The above analysis on the decision-making of the RL agent demonstrates that the proposed algorithm with the state definition $s_{k}=[n_{ij,k},\Delta n_{ij,k},\Delta^{2}n_{ij,k},M_{ij,k}]$ is capable of rejecting misinformation and thus adapting more effectively to disruptions than the baseline agent, leading to antifragile performance as the disruption magnitude intensifies.
@@ 6.2 Performance under real-world limited observability
As discussed in Section 4.1, obtaining real-world measurements of $n_{ij}$ and $q_{ij}$ is highly challenging. Therefore, based on Eq. 8 and 9, we evaluate the performance of the baseline and proposed antifragile RL-based algorithms under conditions of real-world limited observability. It is important to highlight that the performance of the proposed algorithm under limited observability is not directly comparable to MPC or MPC-MHE under full observability, given their fundamentally different observability requirements. Fig. LABEL:fig:_d_obs and LABEL:fig:_d_obs_un illustrate the performance of algorithms in the presence of linearly increasing demand disruptions, modeled either deterministically or with disruption magnitude uncertainties.
@@
In Fig. LABEL:fig:_incr_d_obs, while the TTS curve of the baseline RL algorithm in light coral is approximately the same under both full and limited observability, the performance curve of the antifragile RL algorithm under limited observability in light teal falls between that of the baseline RL and the antifragile RL with full observability. This indicates that the algorithm sacrifices some performance in exchange for depending on more realistic observable measurements. The TTS difference in Fig. LABEL:fig:_incr_d_p_obs marks the performance difference between the baseline and antifragile RL-based algorithms under limited observability, taking an average of $6.1\%$ performance improvement and reaching $15.0\%$ at the end of the simulation. It should be highlighted that the performance trade-off to achieve realistic observability, i.e., the performance difference between limited (light teal) and full observability (blue) of the proposed algorithm, is an average of $4.0\%$. The distribution skewness is quantitatively illustrated in Fig. LABEL:fig:_incr_d_sk_obs. Although taking a higher skewness value at the onset of disruption, the proposed method yields a rather low ultimate skewness of $0.49$ under limited observability, lying between the skewness observed with full observability with $0.46$ and that of the baseline RL under limited observability with $0.62$.
@@
When under demand disruption with magnitude stochasticity as shown in Fig. LABEL:fig:_incr_d_obs_un, the proposed RL algorithm suffers from slightly larger deterioration compared to the baseline RL, mainly because the performance of the baseline RL is close to network gridlock. Still, the TTS difference in Fig. LABEL:fig:_incr_d_p_obs_un showcases the superiority of the proposed algorithm with an average of $4.2\%$ performance gain and reaches $8.2\%$ at the final episode. The performance loss for the trade-off of observability is $4.3\%$. The skewness under disruption stochasticity exhibits some volatility, which shows a value of $0.61$ and is identical to that of the baseline RL under limited observability.
@@
In Fig. LABEL:fig:_s_obs and LABEL:fig:_s_obs_un, the performance of both algorithms under supply disruptions with limited observability follows similar trends to those under demand disruption. The shaded area in Fig. LABEL:fig:_incr_d_p_obs_un exhibits the TTS difference between the two algorithms with limited observability, which yields an average performance gain of $11.1\%$ and achieves a final value of $30.7\%$. The observability trade-off is $5.5\%$. Likewise, its skewness $0.70$ at episode $25$ is also positioned between the baseline RL with limited observability and the antifragile RL with full observability. When under disruption magnitude stochasticity in Fig. LABEL:fig:_incr_s_obs_un, as explained, the performance of the baseline RL method improves at high disruption magnitude, because the network is close to gridlock. Fig. LABEL:fig:_incr_s_p_obs_un exhibits an ultimate performance gain of $12.5\%$. Although this scenario shows the highest observability trade-off of $10.1\%$, the skewness of the proposed antifragile RL is rather low with a value of $0.69$ compared to the other algorithms.
6.3 Simulation study based on real-world data
To demonstrate the applicability of the proposed algorithm in real-world cases and test its robustness, particularly under a time-varying demand profile, we further conduct a numerical simulation using data and models derived from real-world measurements in Fig. 9. Instead of a cordon-shaped network, here we follow the network as in Fulton et al. (2025), where the city center of Zurich is divided into the left and right regions, represented by blue and red links for motorized vehicles, as shown in Fig. 9(a). Such an experimental setup avoids the complexity of through traffic passing by the city center in real-world scenarios. The green dots represent the coordinates of 500 loop detectors spread across the studied region. The measurements from the relevant detectors are used to approximate the MFD for each region. Note that due to the insignificant congestion branch in Zurich, we substituted the third-degree polynomial approximation originally applied in Zhou and Gayah (2021) with the infrastructural potential $\lambda$ method introduced in (Ambühl et al., 2020), which is a semi-analytical method that guarantees the existence of a second root, as illustrated in Fig. 9(b). The switch of the approximation function can also be used to demonstrate the insensitivity of our proposed algorithm to different MFD functional forms. The $\lambda$ method requires certain parameters with physical meanings, including free-flow speed, maximal capacity, backward wave speed, and maximal density, which are also provided in (Ambühl et al., 2020) for the studied region. In addition, by taking the average trip length and total lane length into consideration, the qkq-k MFD is then converted to the mnm-n MFD, as described in Sun et al. (2024).
(a) The city center of Zurich.
(b) Approximated MFDs with loop detector data.
Figure 9: Simulation environment constructed from the real-world network and measurements.
A realistic morning peak hour demand profile in terms of OD pairs $q_{ij,k}$ is also provided in Fulton et al. (2025) for the network described above, which contains strong time-varying stochasticity. It should be noted that real-world OD data is near-impossible to acquire, and the demand profile is calibrated through the methodology described in Ni et al. (2025). Due to the low level of inter-regional traffic flow, to ensure that perimeter control takes effect, a portion of intra-regional traffic within the right region is relocated with its new origin from the left region and destination to the right region. As both the supply and the demand based on real-world datasets in the city center of Zurich have been constructed, to demonstrate the antifragile properties of the proposed algorithm, realistic disruptions need to be approximated through historical data as well. First, the annual average traffic from the relevant loop detector measurements in 2020 is computed as the starting threshold and used as the initial training without disruptions. Then we search for the traffic flow from the beginning of 2021 that exceeds the previous threshold and update the threshold with the new peak. After searching for the whole year, a list of 10 events is summarized and illustrated by the dashed dark red curve with nodes in Fig. 10(a), with the maximal disruption magnitude as $24.5\%$ additional demand compared to the average traffic of the previous year. To tackle the time-varying demand dynamics, a moving window of 10 steps is leveraged to compute the first and second differences in Eq. 8 as well as the redundancy term in Eq. 11. The performance curves from MPC, baseline RL, and the antifragile RL algorithm are presented in Fig. 10(a), where the proposed algorithm not only converges to a lower TTS compared to the baseline algorithm, but also yields an increasingly superior performance with growing disruptions calibrated with real-world measurements. Likewise, Fig. 10(b) illustrates the distribution skewness of the three examined algorithms over the last 10 episodes and showcases that the proposed algorithm offers the lowest skewness in general, indicating its antifragile properties and the applicability under real-world scenarios.
(a) Performance curves in TTS.
(b) Skewness up to the $n$-th episode.
Figure 10: Performance curves and skewness under real-world disruptions.
6.4 Antifragile module on recent RL algorithms
So far, this study builds on the DDPG agent for perimeter control proposed in Zhou and Gayah (2021) and develops an antifragile perimeter control algorithm by augmenting the state space and shaping the reward function of the DDPG agent. The methodology functions as an add-on to the existing DDPG algorithm and enhances its performance under disruptions of previously unseen magnitudes. At the same time, however, recent advances in RL offer various advantages, including improved sample efficiency, training stability, and hyperparameter robustness, among others. Therefore, we apply our proposed methodology to the state-of-the-art algorithms to assess whether, and to what extent, such algorithms would benefit as well from the inclusion of derivatives and reward shaping under increasingly severe disruptive events. In addition to DDPG, we consider Twin Delayed Deep Deterministic policy gradient (TD3) and Soft Actor-Critic (SAC) as baseline algorithms. TD3 employs a twin critics mechanism and takes the minimum Q-value estimate to mitigate overestimation, incorporates target policy smoothing to avoid chasing sharp peaks of Q-value, and delays actor updates so that the policy is trained on steadier value estimates (Fujimoto et al., 2018). SAC is a stochastic, off-policy actor–critic algorithm that maximizes the expected return augmented with an entropy bonus and applies twin Q-networks with automatic temperature tuning to achieve stable and sample-efficient exploration (Haarnoja et al., 2018). Both TD3 and SAC have seen growing applications in the discipline of reliability and risk engineering in recent years, such as in Zhou et al. (2026); Ding et al. (2025); Cao et al. (2025).
(a) Performance curves in TTS.
(b) Normalized over baseline DDPG.
(c) Skewness up to the $n$-th episode.
Figure 11: Performance curves under incremental demand disruptions with advanced RL algorithms.
Fig. 11 displays the performance curve in TTS and the distribution skewness of the studied algorithms. The dashed curves mark the baseline algorithms for DDPG, TD3, and SAC with $s_{k}=[n_{ij,k},q_{ij,k}]$, while the solid curves represent the proposed antifragile variants with $s_{k}=[n_{ij,k},\Delta n_{ij,k},\Delta^{2}n_{ij,k},M_{ij,k}]$ and additional reward shaping terms as in Eq. 4. Clearly in Fig. 11(a), the six performance curves under incremental demand disruptions between episode 507550-75 fall into two regimes, dependent on whether the antifragile module has been leveraged. Fig. 11(b) further gauges the enhanced performance between the baselines and the proposed antifragile algorithms with an improvement varying between $20\%-30\%$ approximately. The distribution skewness is also illustrated in Fig. 11(c). The skewness of the solid curves is consistently lower than the dashed curves, with an average value of $0.57$ compared to $0.69$, demonstrating the antifragile properties of the proposed algorithm under disruptions of increasing magnitudes. Both the superior performance and the lower skewness have showcased the applicability of the proposed antifragile methodology regardless of the specific RL algorithm employed.